OpenStudy (anonymous):

the following function shows the relationship between the selling price (s)and profit p(s) in dollars for a company. p(s)=-20s^2+1,400s-12,000 which statement best describes the intervals where the company profit increases decreases or records a maximum? a.it is least when the selling price is $30 b.it is greatest when the selling price is $30 c.it decrease when the selling price increase from $10 to $35 d. it increase when the selling price increases from $10 to $35

OpenStudy (anonymous):

@primeralph @whpalmer4 @agent0smith @robtobey @e.mccormick @douglaswinslowcooper @Kainui @esshotwired @TheRealMeeeee

OpenStudy (whpalmer4):

Have you tried graphing \(p(s)\) over the range of s = 10 to s = 35?

OpenStudy (whpalmer4):

or even just evaluating it at s = 10, s = 30, s = 35 should give you a clue

OpenStudy (anonymous):

i think the answer is c but im not sure

OpenStudy (whpalmer4):

you could also recognize that this is a parabola and find its vertex which would give you the information to answer

OpenStudy (whpalmer4):

it's good that you aren't sure about c being correct because it is not correct :-)

OpenStudy (anonymous):

oh lol okay

OpenStudy (anonymous):

so it must be d then

OpenStudy (whpalmer4):

no need to guess, you can figure it out and know...

OpenStudy (whpalmer4):

it's a parabola in the form \[y = ax^2+bx+c\]where \[a = -20, ~b = 1400, ~c=-12000\]vertex of a parabola is at \[x = -\frac{b}{2a}\]so our vertex is at \[x = -\frac{1400}{2(-20)} = \frac{1400}{40} = 35\]and because the value of \(a\) is negative, the parabola opens down and the vertex is at the maximum value of \(y\)

OpenStudy (whpalmer4):

Here's a graph:

OpenStudy (whpalmer4):

or if you aren't sure that's the vertex, here's a wider graph with a vertical line at s = 35

OpenStudy (anonymous):

okay i'm for sure it's d now

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

thanks for your help

OpenStudy (anonymous):

@ninjasandtigers do you know the answer to this one

OpenStudy (anonymous):

doesn't he already have the answer? @liorafanny

OpenStudy (anonymous):

ya nvm